OpenStudy (anonymous):

How do I find the inverse function of this one: f(x)=(3x-1)/(x+5)

OpenStudy (e.mccormick):

Same as any invese. Put in y for x and x for f(x), then solve for y.

OpenStudy (e.mccormick):

\(f(x)=\dfrac{3x-1}{x+5} \) becomes \(x=\dfrac{3y-1}{y+5} \) Now solve for y.

hero (hero):

@danielle02, do you know how to finish solving for y?

OpenStudy (anonymous):

@Hero no, I don't.

OpenStudy (yttrium):

Do you know cross multiplication? If you, you can apply it.

OpenStudy (yttrium):

That is actually like this. x(y+5) = 3y - 1 expand the left term and after that, isolate all terms having y to the left and all terms not having y to the right. Then, factor out y to the left and then there it is by equating the equation by y.

hero (hero):

To finish solving for y, do this: Multiply both sides by (y + 5): \[x(y + 5) = 3y + 1\] Distribute \(x\) to \(y + 5\): \[xy + 5x = 3y + 1\] Subtract 3y from both sides; Add 5x to both sides: \[xy - 3y = 1 - 5x\] Factor \(y\) from \(xy - 3y\): \[y(x - 3) = 1 - 5x\] Divide both sides by (x - 3) \[y = \frac{1- 5x}{x -3}\] The last step is to replace \(y\) with \(f^{-1}(x)\) \[f^{-1}(x) = \frac{1- 5x}{x -3}\]

OpenStudy (anonymous):

Ok, I understand now. Thanks for the help.